# DLTPose: 6DoF Pose Estimation From Accurate Dense Surface Point Estimates

### v1

@@ 4.5 Ablations
We performed two experiments, to identify the effect of symmetric keypoints, and the effect of the pseudo-symmetric loss on training. The details are described in the Supplementary Material, Sec. S.5. In summary, they show that the use of the symmetric keypoints (Sec. 3.3) improves both ADD-S and $AR$ scores for symmetric objects over KeyGNet keypoints (Table S.1), and that the pseudo-symmetric loss terms improve the ADD(-S) scores for all LM-O objects (Table S.2).
@@ 5 Conclusion
Future work will focus on methods to further improve the accuracy of surface estimation, which will in turn improve pose estimation accuracy. We will also explore other approaches for generating and dynamically ordering symmetric keypoints during training.
@@
$\begin{bmatrix}-2\bar{x}_{k}&-2\bar{y}_{k}&-2\bar{z}_{k}&1&\left(\bar{x}_{k}^{2}+\bar{y}_{k}^{2}+\bar{z}_{k}^{2}-r^{2}\right)\end{bmatrix}$ (S.4)
$\cdot\begin{bmatrix}\bar{x}\\ \bar{y}\\ \bar{z}\\ \left(\bar{x}^{2}+\bar{y}^{2}+\bar{z}^{2}\right)\\ 1\end{bmatrix}=0$
Finally, we simplify the notation to $|\!|\bar{k}|\!|^{2}=\bar{x}_{k}^{2}+\bar{y}_{k}^{2}+\bar{z}_{k}^{2}$ and $|\!|\bar{p}|\!|^{2}=\bar{x}^{2}+\bar{y}^{2}+\bar{z}^{2}$ and stack a series of $N_{k}$ such rows into matrix $A$ to yield Eq. 2:
$\begin{bmatrix}\!\!-2\bar{x}_{k_{1}}\!\!\!\!\!\!\!\!&\!\!\!\!-2\bar{y}_{k_{1}}\!\!\!\!\!\!\!\!&\!\!\!\!\!\!\!\!-2\bar{z}_{k_{1}}\!\!\!\!\!\!\!\!&\!\!\!\!1&\!\!\!\!(|\!|\bar{k}_{1}|\!|^{2}\!-\!\hat{r}_{1}^{2})\\ \vdots&\vdots&\vdots&\!\!\!\!\vdots&\vdots\\ \!-2\bar{x}_{k_{N_{k}}}\!\!\!\!\!\!\!\!&-2\bar{y}_{k_{N_{k}}}\!\!\!\!\!\!\!\!&-2\bar{z}_{k_{N_{k}}}\!\!\!\!&\!\!\!\!1&\!\!\!\!(|\!|\bar{k}_{N_{k}}|\!|^{2}\!-\!\hat{r}_{N_{k}}^{2})\\ \end{bmatrix}\begin{bmatrix}\bar{x}\\ \bar{y}\\ \bar{z}\\ |\!|\bar{p}|\!|^{2}\\ 1\end{bmatrix}=0.$
@@
17:       Compute normal vector: $\vec{n}_{i}=\frac{\vec{e}_{1}\times\vec{e}_{2}}{||\vec{e}_{1}\times\vec{e}_{2}||}$
@@
Table S.1: Comparison of ADD-S (without and with ICP) and $AR$ metrics for two symmetric objects, using keypoints selected using either KeyGNet [wu2024keygnet] or our symmetric keypoint generation method (Sec. 3.3).
@@
Object w/o ICP w ICP
$\mathcal{L}_{R}$ $\mathcal{L}_{R}+\mathcal{L}_{C}$ $\cal{L}_{\text{total}}$ $\mathcal{L}_{R}$ $\mathcal{L}_{R}+\mathcal{L}_{C}$ $\cal{L}_{\text{total}}$
ape 73.5 73.3 74.2 85.5 86.5 87.1
@@
The detailed ADD(-S) results for the LM and LM-O datasets, along with the AUC results for ADD-S and ADD(-S) on the YCB-V dataset, are provided in Tables S.3S.5. Additionally, Figure S.2– S.4 presents qualitative examples of successful pose recoveries, where red dots represent projected surface points from ground truth poses, while blue dots correspond to those from the estimated poses.
@@
Table S.4 highlights the challenges posed by the LM-O dataset, where occlusions and imperfect object meshes introduce significant difficulties to pose estimation. Despite these challenges, our method achieves the highest mean accuracy of 90.4%, outperforming prior state-of-the-art approaches and showcasing strong robustness in occluded scenarios.
For the YCB-V dataset, Table S.5 reports AUC for both ADD-S and ADD(-S). Without ICP, our approach achieves a mean AUC of 98.3% for ADD-S and 93.9% for ADD(-S). With ICP, the performance further improves to 99.7% for ADD-S and 97.1% for ADD(-S), marking a significant improvement over prior state-of-the-art methods, i.e. +2.5% for ADD-S, and +4.8% for ADD(-S).
@@
Object Dense- Fusion  [wang2019densefusion] RCVPose [Wu] PVN3D [He] REDE [Hua2020REDEEO] Masked- Fusion [pereira2020maskedfusion] PR-GCN [Zhou2021PRGCNAD] EANet [ea-net] IRPE [Jin2024IRPEIR] DLTPose w/o ICP DLTPose
ape 92.3 99.6 97.3 95.6 92.2 99.2 95.1 95.7 99.7 100.0
benchvise 93.2 99.7 99.7 99.4 98.4 99.8 97.5 99.9 99.8 100.0
camera 94.4 99.7 99.7 99.6 98.0 100.0 98.5 97.2 98.7 100.0
can 93.1 99.3 98.0 97.8 97.4 99.3 97.7 100.0 100.0 100.0
cat 96.5 99.7 99.9 99.5 97.7 99.8 97.7 99.6 99.8 100.0
@@
duck 92.3 99.7 99.4 98.0 94.0 98.2 97.3 97.9 100.0 100.0
eggbox 99.8 99.3 99.3 98.6 99.6 99.9 99.6 99.7 99.8 100.0
glue 100.0 99.7 100.0 100.0 100.0 100.0 99.9 99.7 100.0 100.0
holepuncher 92.1 100.0 99.3 98.6 97.3 99.0 96.8 98.3 99.8 99.9
iron 97.0 99.9 99.7 99.8 97.1 100.0 99.4 98.3 98.5 99.0
lamp 95.3 99.5 99.5 99.3 99.0 100.0 99.2 98.2 93.8 100.0
phone 92.8 99.7 99.5 99.5 98.8 99.9 98.7 98.2 97.8 100.0
@@
Object PoseCNN [xiang2018posecnn] PVN3D [He] RCVPose [Wu] REDE [Hua2020REDEEO] PR-GCN [Zhou2021PRGCNAD] ZebraPose [su2022zebrapose] IRPE [Jin2024IRPEIR] DLTPose w/o ICP DLTPose
ape 76.2 33.9 61.3 53.1 40.2 60.4 69.8 74.2 87.1
can 87.4 88.6 93.0 88.5 76.2 95.0 95.4 93.3 96.8
cat 52.2 39.1 51.2 35.9 57.0 62.1 72.4 59.8 71.0
driller 90.3 78.4 78.8 77.8 82.3 94.8 93.0 90.6 97.3
duck 77.7 41.9 53.4 46.2 30.0 64.5 81.1 52.7 84.1
eggbox* 72.2 80.9 82.3 71.8 68.2 73.8 85.3 92.0 98.6
glue* 76.7 68.1 72.9 75.0 67.0 88.7 88.4 89.1 91.1
holepuncher 91.4 74.7 75.8 75.5 97.2 88.4 97.9 91.0 97.5
Mean 78.0 63.2 71.1 65.4 64.8 78.5 85.4 80.3 90.4
@@
Object PoseCNN [xiang2018posecnn] PVN3D [He] ZebraPose [su2022zebrapose] EANet  [ea-net] RCVPose [Wu] IRPE [Jin2024IRPEIR] DLTPose w/o ICP DLTPose
-S (-S) -S (-S) -S (-S) -S (-S) -S (-S) -S (-S) -S (-S) -S (-S)
002_master_chef_can 95.8 69.0 95.2 79.3 96.3 80.0 96.6 96.2 100.0 100.0 99.7 96.5 100.0 97.5
003_cracker_box 91.8 80.7 94.4 91.5 93.5 88.6 96.2 97.9 99.0 96.8 99.8 98.8 99.8 99.8
004_sugar_box 98.2 97.2 97.9 96.9 95.5 91.6 97.8 97.9 96.8 95.0 99.3 95.0 99.9 99.8
005_tomato_soup_can 94.5 94.3 95.9 89.0 94.4 90.3 96.1 99.0 98.2 94.1 100.0 99.6 100.0 99.8
006_mustard_bottle 98.4 87.0 98.3 97.9 96.1 93.0 96.9 98.2 99.0 95.3 100.0 99.9 100.0 100.0
007_tuna_fish_can 98.4 97.9 96.7 90.7 97.7 94.8 97.1 98.6 98.0 98.0 98.8 90.9 99.1 92.3
008_pudding_box 97.9 96.6 98.2 97.1 94.2 84.4 94.8 98.1 100.0 100.0 100.0 99.8 100.0 100.0
009_gelatin_box 98.8 96.6 98.8 98.3 97.8 88.4 97.1 98.4 99.0 98.9 99.8 97.0 100.0 100.0
010_potted_meat_can 92.8 83.8 93.8 87.9 93.8 84.0 97.2 98.4 86.0 74.4 99.1 97.4 99.7 98.1
011_banana 96.9 92.6 98.2 96.0 92.3 84.3 97.1 98.3 99.9 98.9 96.9 89.9 99.9 97.1
019_pitcher_base 97.8 92.3 97.6 96.9 89.8 89.0 98.0 97.2 91.2 87.8 98.3 84.7 99.5 87.9
021_bleach_cleanser 96.8 92.3 97.2 95.9 89.8 89.0 98.0 99.6 79.3 74.0 98.8 90.3 99.8 96.9
@@
037_scissors 92.2 84.8 92.4 95.0 87.2 91.9 94.0 96.4 99.7 99.7 80.1 67.4 97.1 87.0
040_large_marker 97.2 47.3 98.1 91.6 97.6 89.7 94.0 96.6 97.6 89.7 99.9 95.2 99.9 96.1

### v2

@@ 4.4 Ablations
We conducted additional experiments to evaluate: (i) the impact of symmetric keypoints, (ii) the effect of incorporating the pseudo-symmetric loss during training, (iii) the performance difference between per-object models and a single model trained on all objects, (iv) the influence of using object segmentations versus bounding boxes during training, and (v) the effect of varying the number of keypoints (4, 8, 12, 16) on object pose estimation accuracy. The details are described in the Supplementary Material, Sec. S.5. In summary, symmetric keypoints significantly enhance ADD-S and AR metrics by explicitly addressing pose ambiguities for symmetric objects. Incorporating pseudo-symmetric loss improves ADD(-S) scores across LM-O objects by refining radial map regression accuracy. Per-object models achieve higher AR compared to a unified model, as they effectively learn object-specific geometric constraints. Employing fewer keypoints (especially 4) achieves superior or comparable accuracy by reducing per-channel radial regression errors. Lastly, segmented inputs consistently outperform bounding boxes due to more accurate localization of object boundaries, yielding precise radial predictions.
@@ 5 Conclusion
A limitation of DLTPose is that the DLT formulation requires non-coplanar keypoints, and when this condition is not met, the resulting pose estimates become unstable. A further limitation is that the approach employs distinct keypoint strategies for symmetric and asymmetric objects, with symmetric keypoints manually defined.
Future work will focus on methods to further improve the accuracy of surface estimation, which will in turn improve pose estimation accuracy. We will also explore other approaches for automatically generating and dynamically ordering symmetric keypoints during training.
@@
$\begin{bmatrix}-2\bar{x}_{k}&-2\bar{y}_{k}&-2\bar{z}_{k}&1&(\bar{x}_{k}^{2}+\bar{y}_{k}^{2}+\bar{z}_{k}^{2}-r^{2})\end{bmatrix}\\ \hskip 99.58464pt\begin{bmatrix}\bar{x}\\ \bar{y}\\ \bar{z}\\ (\bar{x}^{2}+\bar{y}^{2}+\bar{z}^{2})\\ 1\end{bmatrix}=0.\;\;\;$ (S.4)
Finally, we simplify the notation to $\bar{k}|\!|^{2}=\bar{x}_{k}^{2}+\bar{y}_{k}^{2}+\bar{z}_{k}^{2}$ and $\bar{p}|\!|^{2}=\bar{x}^{2}+\bar{y}^{2}+\bar{z}^{2}$ and stack a series of $N_{k}$ such rows into matrix $A$ to yield Eq. 2:
[…]
@@
[…]
@@
Figure S.1: Average recall components (MSSD, MSPD, and ARAR) vs. standard deviation of injected noise on the estimated LM-O object surface points, both with (solid lines) and without (dotted lines) ICP refinement.
@@
Table S.1: Comparison of ADD-S (without and with ICP) and ARAR metrics for three symmetric objects, using keypoints selected using either KeyGNet [wu2024keygnet] or our symmetric keypoint generation method (Sec. 3.3).
@@
Loss w/o ICP w ICP
$\mathcal{L}_{R}$ (Eq. 3)
$\mathcal{L}_{C}$ (Eq. 4)
$\mathcal{L}_{P}$ (Eq. 5)
Objects ape 73.5 73.3 74.2 85.5 86.5 87.1
@@
Effect of Training a Single Model on the Entire LM-O Dataset. In this ablation study, we investigate the consequences of training one unified model on the LM-O dataset, which contains both symmetric and asymmetric objects. We optimize the model using the total loss $\mathcal{L}_{\text{total}}$ (Eq. 6) and assess its performance via the ADD-S and AR metrics. Table S.3 reports the ADD(-S) scores on LM-O for multiple objects, both before and after ICP refinement. Training a single model on the full dataset yields a 0.2 % increase in mean ADD(-S) without ICP but the same ADD(-S) with ICP.
Table S.4 reports the Average Recall scores on LM-O for multiple objects after ICP refinement. Notably, the Eggbox object trained with symmetric keypoints, shows a significant drop in AR under the single model setup (0.579) compared to training a separate model per object (0.691). This drop is caused by a mismatch in training objectives: the single model is trained with the total loss $\mathcal{L}_{\text{total}}$ (Eq. 6), whereas symmetric objects are more effectively trained using only the radial loss $\mathcal{L}_{R}$ (Eq. 3). As a result, the shared model fails to preserve the geometric constraints specific to symmetric object. Overall, the mean AR for the single model approach on LM-O is 0.784, which is 1.6% lower than the 0.800 mean AR achieved when training individual models per object.
Object One Model per Object One Model per Dataset
w/o ICP w ICP w/o ICP w ICP
ape 74.2 87.1 77.2 86.5
can 93.3 96.8 95.3 98.2
cat 59.8 71.0 63.1 71.8
driller 90.6 97.3 87.6 96.1
duck 79.1 88.5 77.2 88.0
[…]
holepuncher 90.1 97.5 95.1 98.6
Mean 83.6 91.0 83.8 91.0
Table S.3: Comparison of ADD(-S) metric, on One Model per Object and One Model per dataset, on LM-O for models trained with KeyGNet and symmetric keypoints using the total loss $\mathcal{L}_{\text{total}}$ (Eq. 6).
Object One Model per Object One Model per Dataset
$\textit{AR}_{VSD}$ $\textit{AR}_{MSSD}$ $\textit{AR}_{MSPD}$ Mean AR $\textit{AR}_{VSD}$ $\textit{AR}_{MSSD}$ $\textit{AR}_{MSPD}$ Mean AR
ape 0.638 0.830 0.912 0.795 0.645 0.822 0.906 0.791
can 0.737 0.914 0.921 0.857 0.744 0.921 0.926 0.864
cat 0.568 0.713 0.797 0.693 0.552 0.697 0.788 0.679
driller 0.788 0.937 0.919 0.881 0.773 0.922 0.901 0.865
duck 0.725 0.808 0.903 0.812 0.728 0.812 0.896 0.812
[…]
holepuncher 0.738 0.901 0.932 0.859 0.741 0.908 0.930 0.859
Table S.4: Comparison of Average Recall metric per object after refinement, on One Model per Object and One Model per dataset, on LM-O for models trained with KeyGNet and symmetric keypoints using the total loss $\mathcal{L}_{\text{total}}$ (Eq. 6).
Effect of Training models with segmented objects. In this ablation study, we evaluate the effect of using segmentation masks versus object bounding boxes as input regions for radial map regression. The model is trained separately for objects in the LM-O dataset, using the total loss $\mathcal{L}{\text{total}}$ (Eq. 6) for asymmetric objects. Performance is assessed using the ADD-S and AR metrics.
Object Segmented Object ROI Object Bounding Box ROI
w/o ICP w ICP w/o ICP w ICP
ape 74.2 87.1 65.7 83.9
can 93.3 96.8 90.3 95.9
cat 59.8 71.0 45.8 66.9
driller 90.6 97.3 85.9 95.6
Table S.5: Comparison of ADD(-S) scores on the LM-O dataset for models trained with KeyGNet and symmetric keypoints, using segmented object regions and object bounding boxes as input ROIs.
Object Segmented Object ROI Object Bounding Box ROI
$\textit{AR}_{VSD}$ $\textit{AR}_{MSSD}$ $\textit{AR}_{MSPD}$ Mean AR $\textit{AR}_{VSD}$ $\textit{AR}_{MSSD}$ $\textit{AR}_{MSPD}$ Mean AR
ape 0.638 0.830 0.912 0.795 0.607 0.789 0.891 0.763
can 0.737 0.914 0.921 0.857 0.726 0.895 0.898 0.834
cat 0.568 0.713 0.797 0.693 0.525 0.667 0.768 0.654
driller 0.788 0.937 0.919 0.881 0.764 0.922 0.901 0.862
Table S.6: Comparison of per-object Average Recall after refinement on the LM-O dataset for models trained with KeyGNet and symmetric keypoints, using segmented object regions and object bounding boxes as input ROIs.
On segmented object ROIs, the model demonstrates improved radial map regression, which in turn enhances surface representation and overall pose estimation accuracy. As shown in Table S.5, the ADD(-S) scores are consistently higher across the objects when using segmented regions compared to bounding box inputs. Similarly, Table S.6 shows that Average Recall metrics, including $\textit{AR}_{VSD}$, $\textit{AR}_{MSSD}$, and $\textit{AR}_{MSPD}$, are superior for each object when the model is trained with segmented object ROIs. These results highlight the benefit of providing the model with more precise object boundaries during training, enabling it to learn more accurate radial offsets.
Effect of Number of Keypoints in DLT Estimation. In this ablation study, we evaluate how varying the number of keypoints influences the Direct Linear Transform (DLT) method for object surface and pose estimation. The models are individually trained on a few objects from the LM-O dataset, incrementally using 4, 8, 12, and 16 keypoints. Each model is assessed based on its performance in predicting accurate object poses, employing the ADD-S and Average Recall metric. As shown in Tables S.7 and S.8, models trained with only 4 keypoints consistently achieved either the best or second-best performance. We attribute this to reduced per-channel error in the radial regression task when using fewer keypoints. As the number of keypoints increases, the likelihood of noisy or inconsistent radial predictions across channels grows, potentially degrading overall accuracy. Therefore, the main results reported in this paper are derived using the 4-keypoint configuration.
Object 4 Keypoints 8 Keypoints 12 Keypoints 16 Keypoints
w/o ICP w ICP w/o ICP w ICP w/o ICP w ICP w/o ICP w ICP
ape 74.2 87.1 73.0 85.0 71.6 83.6 74.1 85.0
can 93.3 96.8 90.4 96.5 95.9 98.5 90.3 97.1
cat 59.8 71.0 59.3 70.4 59.4 70.2 55.4 68.9
driller 90.6 97.3 86.2 94.2 89.0 95.4 87.6 95.9
Table S.7: ADD-S of four LM-O objects with DLTPose as the number of KeyGNet keypoints increases from 4 to 16.
Object 4 Keypoints 8 Keypoints 12 Keypoints 16 Keypoints
ape 79.5 77.6 76.2 77.1
can 85.7 84.0 87.9 84.7
cat 69.3 69.9 68.5 68.1
driller 88.1 86.9 87.7 88.5
Table S.8: Average Recall of four LM-O objects with DLTPose (w ICP) as the number of KeyGNet keypoints increases from 4 to 16.
@@
The detailed ADD(-S) results for the LM and LM-O datasets, along with the AUC results for ADD-S and ADD(-S) on the YCB-V dataset, are provided in Tables S.9S.11. Additionally, Figure S.3– S.5 presents qualitative examples of successful pose recoveries, where red dots represent projected surface points from ground truth poses, while blue dots correspond to those from the estimated poses.
@@
Table S.10 highlights the challenges posed by the LM-O dataset, where occlusions and imperfect object meshes introduce significant difficulties to pose estimation. Despite these challenges, our method achieves the highest mean accuracy of 91.0%, outperforming prior state-of-the-art approaches and showcasing strong robustness in occluded scenarios.
For the YCB-V dataset, Table S.11 reports AUC for both ADD-S and ADD(-S). Without ICP, our approach achieves a mean AUC of 98.3% for ADD-S and 94.3% for ADD(-S). With ICP, the performance further improves to 99.7% for ADD-S and 97.4% for ADD(-S), marking a significant improvement over prior state-of-the-art methods.
@@
Object
Dense-Fusion [wang2019densefusion] RCVPose [Wu] PVN3D [He] REDE [Hua2020REDEEO] Masked-Fusion [pereira2020maskedfusion] PR-GCN [Zhou2021PRGCNAD] EANet [ea-net] IRPE [Jin2024IRPEIR] DLTPose w/o ICP DLTPose
ape 92.3 99.6 97.3 95.6 92.2 99.2 95.1 95.7 99.7 100
benchvise 93.2 99.7 99.7 99.4 98.4 99.8 97.5 99.9 99.8 100
camera 94.4 99.7 99.7 99.6 98.0 100 98.5 97.2 98.7 100
can 93.1 99.3 98.0 97.8 97.4 99.3 97.7 100 100 100
cat 96.5 99.7 99.9 99.5 97.7 99.8 97.7 99.6 99.8 100
@@
duck 92.3 99.7 99.4 98.0 94.0 98.2 97.3 97.9 100 100
eggbox 99.8 99.3 99.3 98.6 99.6 99.9 99.6 99.7 99.8 100
glue 100 99.7 100 100 100 100 99.9 99.7 100 100
holepuncher 92.1 100 99.3 98.6 97.3 99.0 96.8 98.3 99.8 99.9
iron 97.0 99.9 99.7 99.8 97.1 100 99.4 98.3 98.5 99.0
lamp 95.3 99.5 99.5 99.3 99.0 100 99.2 98.2 93.8 100
phone 92.8 99.7 99.5 99.5 98.8 99.9 98.7 98.2 97.8 100
@@
Object
PoseCNN [xiang2018posecnn] PVN3D [He] RCVPose [Wu] REDE [Hua2020REDEEO] PR-GCN [Zhou2021PRGCNAD] ZebraPose [su2022zebrapose] IRPE [Jin2024IRPEIR] EPRO-GDR [10896093] DFTr [zhou2023deep] HiPose [lin2024hipose] DLTPose w/o ICP DLTPose
ape 76.2 33.9 61.3 53.1 40.2 60.4 69.8 64.1 76.6 78.0 74.2 87.1
can 87.4 88.6 93.0 88.5 76.2 95.0 95.4 98.5 96.1 98.9 93.3 96.8
cat 52.2 39.1 51.2 35.9 57.0 62.1 72.4 88.9 52.2 87.5 59.8 71.0
driller 90.3 78.4 78.8 77.8 82.3 94.8 93.0 96.5 95.8 97.8 90.6 97.3
duck 77.7 41.9 53.4 46.2 30.0 64.5 81.1 82.8 72.3 85.3 79.1 88.5
eggbox* 72.2 80.9 82.3 71.8 68.2 73.8 85.3 74.4 75.3 80.3 92.0 98.6
glue* 76.7 68.1 72.9 75.0 67.0 88.7 88.4 95.0 79.3 94.1 89.1 91.1
holepuncher 91.4 74.7 75.8 75.5 97.2 88.4 97.9 96.5 86.8 95.2 91.0 97.5
Mean 78.0 63.2 71.1 65.4 64.8 78.5 85.4 88.7 77.7 89.6 83.6 91.0
@@
Object PoseCNN [xiang2018posecnn] PVN3D [He] ZebraPose [su2022zebrapose] EANet  [ea-net] RCVPose [Wu] DFTr [zhou2023deep] HiPose [lin2024hipose] IRPE [Jin2024IRPEIR] DLTPose w/o ICP DLTPose
-S (-S) -S (-S) -S (-S) -S (-S) -S (-S) -S (-S) -S (-S) -S (-S) -S (-S) -S (-S)
002_master_chef_can 95.8 69.0 95.2 79.3 96.3 80.0 96.6 96.2 97.0 92.3 96.4 86.2 100 100 99.7 96.5 100 97.5
003_cracker_box 91.8 80.7 94.4 91.5 93.5 88.6 96.2 97.9 95.9 93.9 97.7 96.7 99.0 96.8 99.8 98.8 99.8 99.8
004_sugar_box 98.2 97.2 97.9 96.9 95.5 91.6 97.8 97.9 97.1 95.5 98.2 97.1 96.8 95.0 99.3 95.0 99.9 99.8
005_tomato_soup_can 94.5 94.3 95.9 89.0 94.4 90.3 96.1 99.0 95.6 92.6 97.0 95.1 98.2 94.1 100 99.6 100 99.8
006_mustard_bottle 98.4 87.0 98.3 97.9 96.1 93.0 96.9 98.2 97.6 96.3 98.4 96.9 99.0 95.3 100 99.9 100 100
007_tuna_fish_can 98.4 97.9 96.7 90.7 97.7 94.8 97.1 98.6 97.3 94.5 97.8 96.2 98.0 98.0 99.1 98.0 99.3 98.2
008_pudding_box 97.9 96.6 98.2 97.1 94.2 84.4 94.8 98.1 97.4 95.7 98.8 98.1 100 100 100 99.8 100 100
009_gelatin_box 98.8 96.6 98.8 98.3 97.8 88.4 97.1 98.4 97.6 96.3 98.9 97.8 99.0 98.9 99.8 97.0 100 100
010_potted_meat_can 92.8 83.8 93.8 87.9 93.8 84.0 97.2 98.4 95.9 92.1 93.5 83.4 86.0 74.4 99.1 97.4 99.7 98.1
011_banana 96.9 92.6 98.2 96.0 92.3 84.3 97.1 98.3 97.1 95.0 98.6 96.3 99.9 98.9 96.9 89.9 99.9 97.1
019_pitcher_base 97.8 92.3 97.6 96.9 89.8 89.0 98.0 97.2 96.0 93.1 96.8 93.2 91.2 87.8 98.3 84.7 99.5 87.9
021_bleach_cleanser 96.8 92.3 97.2 95.9 89.8 89.0 98.0 99.6 96.8 94.9 97.1 94.0 79.3 74.0 98.8 90.3 99.8 96.9
@@
037_scissors 92.2 84.8 92.4 95.0 87.2 91.9 94.0 96.4 97.2 93.3 98.3 96.8 99.7 99.7 97.3 87.5 99.1 92.1
040_large_marker 97.2 47.3 98.1 91.6 97.6 89.7 94.0 96.6 96.9 92.7 98.6 94.3 97.6 89.7 99.9 95.2 99.9 96.1